OpenStudy (anonymous):

Find the interest earned on an investment with a principal = $880, interest rate = 3 % per year and time = 9 months. $19.80 $899.80 $237.60 $2,376

hartnn (hartnn):

Use the formula I = P*N*R where I = interest P = principal amount = 880 here R = rate of interest = 3% = 3/100 here N = time in years = 9/12 here :)

OpenStudy (anonymous):

i guess we have to assume it is compounded monthly for this problem although it doesnt say that

hartnn (hartnn):

ohhhh...but when i used this formula, i got an answer which is in the choices, so i didn't try compound interest